# Tokenization Matters! Degrading Large Language Models through Challenging Their Tokenization

### v1

@@ 3.3 ADT-Auto generation
Given the inefficiency of constructing dataset manually, we further develop an automatic construction framework for the dataset to challenge LLMs’ tokenization.
As discussed in Section 3.2, the challenges of tokenization in English are less severe than in Chinese. Consequently, we primarily concentrate on the automatic generation of dataset in Chinese. Figure 3 illustrates the process of automatically constructing our dataset ADT-Auto.
@@ 3.3.1 Word-pair matching
The automatic generation of dataset is also based on the exported vocabularies. From the vocabularies, we first seek some qualified word pairs. Given two words Word 1 and Word 2, they are considered to be match when meeting the following criteria: The suffix (or whole) of Word 1 can be concatenated with the prefix (or whole) of Word 2, as a token (word) existing in the vocabulary, denoted as Trap Word. Meanwhile, the concatenation of Word 1 and Word 2 should not be a token existing in the vocabulary.
@@
Above clarifications indicate that the concatenation of Word 1 and Word 2 (as ‘ABCD’ in Figure 3) corresponds to the challenging span in the construction process of ADT-Human. Our objective is to compose a sentence that can not only convey the semantic essence of both Word 1 and Word 2, but also induce LLM to isolate Trap Word during the tokenization process. Notably, we neglect the situation that the concatenation of Word 1 and Word 2 is just Trap Word, because it would not pose challenges to LLMs’ tokenization.
To augment the efficacy of Trap Word upon matching, we also consider the word-pair matching criterion that the remaining parts of Word 1 and Word 2 excluded by Trap Word should also exist in the vocabulary as a token (word). Furthermore, the first or last character of Trap Word cannot be a Chinese stop character.
@@ 3.3.2 Instance generalization
With each Word 1, Word 2 and their corresponding Trap Word, we harness GPT-4 to generate an instance of ADT-Auto, which consists of a sentence and a corresponding question. Specifically, we prompt GPT-4 to generate a sentence that is used to challenge the tokenization of LLMs. To this end, the generated sentence is required to include the concatenation of Word 1 and Word 2 (inevitably including Trap Word). In addition, GPT-4 is also required to devise a question related to the generated sentence, which is used to evaluate LLMs’ tokenization performance through their answers. The instance must meet the criteria as below: The sentence should convey the meanings of both Word 1 and Word 2, while the question’s answer should come from Word 1, Word 2, Trap Word or their combination. Thus, the influence of LLMs’ incorrect tokenization can be identified by checking their answers to the question.
The prompt for GPT-4 to generate ADT-Auto instances also includes some demonstration examples in addition to the task instructions, which is illustrated in Appendix B.
@@ 3.3.3 Filtering based on Trap Word
The goal of our dataset is to expose the flaw of LLMs’ tokenization. Therefore, we will only retain the instances that can induce the tokenization problem of LLMs. To this end, we check each instance generated at the previous step. For these instances, the presence of Trap Word implies a challenging case that is likely to induce tokenization problems. Accordingly, we first select an open-source LLM and get the Trap Word set based on its vocabulary. Given an instance, we retain it if a certain Trap Word is both in the LLM’s tokenization list for the instance’s sentence (as List 1 in Figure 3) and in its tokenization list for the answer (as List 2 in Figure 3). Such filtering criterion indicates that the LLM commits tokenization errors on understanding the sentence and response for the instance. Thus, this instance is also likely to challenge other LLMs’ tokenization.
@@ 3.3.4 Manual filtering
To ensure that the retained instances can induce tokenization problems and meanwhile have reasonable expressions, we further adopt the manual assessment for instances. Specifically, we further select the high-qualify instances with considering their sentence expressions and LLMs’ responses. Notably, we might still retain some instances to which the used LLM has correct response, since the other LLMs are still likely to commit inaccurate tokenization for these instances, resulting in unsatisfactory responses.
Due to space limitation, we take Qwen-7B as an example to illustrate the process of generating ADT-Auto instances. There are 24,953 Chinese tokens in Qwen-7B’s vocabulary, and after word-pair matching, 1,764,692 word-pairs are obtained. From the matching word-pairs, 8,000 pairs are selected randomly and then used for instance generation by GPT-4. Due to the inherent stochastic characteristic of LLMs on response generation, we conduct three iterations of filtering operations introduced in Subsection 3.3.3 to get more qualified instances, and thus retain 894 instances. Next, through the manual filtering introduced in Subsection 3.3.4, we retain 231 for ADT-Auto in the end.
@@ 4.1 Experiment setup
Considering the open-source LLMs used in our experiments, we select the LLMs previously used in the construction of ADT-Human, including Chatglm3-6B, Baichuan2-13B-Chat, Yi-34B-Chat, Qwen-7B-Chat, and Qwen1.5-72B-Chat for Chinese data, as well as Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1 for English data. We test these LLMs using both locally deployed versions and API versions. For the closed-source LLMs used in our experiments, we test GPT-4o, GPT-4, GPT-3.5-Turbo, Qwen2.5-max, step-1-8k [^5] https://platform.stepfun.com/docs/llm/text, moonshot-v1-8k [^6] https://platform.moonshot.cn/docs, ERNIE-3.5-8K777 https://cloud.baidu.com/doc/WENXINWORKSHOP/s/jlil56u11 for Chinese data, and GPT-4o, GPT-4 and GPT-3.5-Turbo for English data.
@@ 4.2 ADT-Human’s challenges to LLMs
Firstly, we investigate the challenges posed by our manually constructed dataset ADT-Human to LLMs. To this end, we evaluate the experimented LLMs’ performance on the dataset, through counting the number of incorrect answers generated by the LLMs for the questions in the instances. Recalling the process of manually composing challenging spans introduced in Section 3.2, the span ‘BC’ in Table 2 is in fact a Trap Word mentioned in Section 3.3. Hence, for a given LLM, its answer including a Trap Word is identified as its inaccurate response for the question undoubtedly. Such a case also indicates that the LLM generates a wrong token segmentation.
The percentage of incorrect responses provided by the LLMs for the Chinese data (corresponding to the four LLMs’ vocabularies) and English data (corresponding to the two LLMs’ vocabularies) are presented in Table 3 and Table 4, respectively. It should be noted that there is no API version of Chatglm3-6B. The results shows that the ADT-Human poses a significant challenge to both open-source and closed-source LLMs on their tokenization, resulting in very high rates of inaccurate responses.
@@
Model Source LLM of vocabulary Overall error rate Chatglm3 Baichuan2 Yi Qwen Open-source (Local) Chatglm3-6B 100.00 100.00 100.00 90.91 97.73 Baichuan2-13B-Chat 90.91 100.00 81.82 100.00 93.18 Yi-34B-Chat 72.73 63.64 100.00 100.00 84.09 Qwen-7B-Chat 100.00 72.73 90.91 100.00 90.91 Qwen1.5-72B-Chat 90.91 45.45 81.82 100.00 79.55 Open-source (API) Baichuan2-13B-Chat 100.00 100.00 90.91 100.00 97.73 Yi-34B-Chat 81.82 54.55 100.00 90.91 81.82 Qwen-7B-Chat 100.00 81.82 81.82 100.00 90.91 Qwen1.5-72B-Chat 90.91 54.55 81.82 100.00 81.82 Closed-source GPT-4o 72.73 27.27 54.55 90.91 61.36 GPT-4 81.82 45.45 45.45 27.28 50.00 GPT-3.5-Turbo 72.73 27.27 72.73 72.73 61.36 Qwen2.5-max 90.91 72.73 90.91 100.00 88.64 step-1-8k 63.64 18.18 72.73 63.64 54.55 moonshot-v1-8k 81.82 27.27 100.00 81.82 72.73 ERNIE-3.5-8K 72.73 54.55 54.55 72.73 63.64
@@
It is also worth mentioning that the recent state-of-the-art (SOTA) GPT-4o cannot outperform GPT-4 (higher error rate), implying that the advancements of these LLMs have not yet addressed this primary but challenging problem.
@@ 4.3 ADT-Auto’s challenges to LLMs
Similar to the investigation of ADT-Human’s challenges to LLMs’ tokenization, we also evaluate the LLMs’ performance on ADT-Auto. We still test five open-source LLMs with their both locally deployed and API versions, and seven closed-source LLMs only having APIs.
@@
Table 5: Error rates of answers on ADT-Auto.
Model Fraction Error rate(%)
Open-source (Local) Chatglm3-6B 156/231 67.53
Baichuan2-13B-Chat 147/231 63.64
Yi-34B-Chat 90/231 38.96
Qwen-7B-Chat 185/231 80.09
Qwen1.5-72B-Chat 93/231 40.26
Open-source (API) Baichuan2-13B-Chat 167/231 72.29
Yi-34B-Chat 80/231 34.63
Qwen-7B-Chat 160/231 69.26
Qwen1.5-72B-Chat 97/231 41.99
Closed-source GPT-4o 75/231 32.47
GPT-4 89/231 38.53
GPT-3.5-Turbo 99/231 42.86
Qwen2.5-max 75/231 32.47
step-1-8k 60/231 25.97
moonshot-v1-8k 60/231 25.97
ERNIE-3.5-8K 51/231 22.08
@@
    Compared with the closed-source LLMs, the open-source LLMs suffer from ADT-Auto’s challenges more apparently. It implies that these closed-source LLMs, as the profit-making flagships of their creator companies, naturally have stronger capabilities than the open-source LLMs that are created for public usage.
@@
    Qwen1.5-72B-Chat has much lower error rates than Qwen-7B-Chat, although they have the same vocabulary. We specially check the instances to which Qwen1.5-72B-Chat gives the correct answers but Qwen-7B-Chat gives the wrong answers, and find the two models have the same incorrect tokenization lists for these instances. It suggests that their different performance is not caused by their tokenization. These results also imply that in the case of incorrect tokenization, the bigger LLMs are more robust and likely to generate correct responses than the smaller LLMs thanks to their stronger capabilities brought by the larger scale.
5 Perspectives and conclusion
Based on our findings in this paper, for the work dedicating to improve LLMs’ tokenization, and so as to improve LLMs’ performance on varied tasks, it is reasonable and promising to expand LLMs’ vocabulary sizes or developing more powerful tokenizers or effective tokenization algorithms. As the idea proposed by Kudo kudo2018subword , employing multiple subword segmentation methods might be a feasible strategy to enhance the robustness of LLMs’ tokenization process. We hope that our work in this paper and the dataset could shed light on the subsequent research on improving LLMs’ tokenization and then the performance on other tasks.
In this paper, we dedicate to deeply investigating the relationship between LLMs’ vulnerability on tokenization and their unsatisfactory responses for some tasks. To this end, we construct an adversarial dataset ADT (Adversarial Dataset for Tokenizer) consisting of a manual constructed subset ADT-Human and an automatically generated subset ADT-Auto, of which each instance includes a sentence and a corresponding question. Our experiments obviously demonstrate that our dataset does challenge the studied open-source and closed-source LLMs’ token segmentation, resulting in their incorrect answers to the instances’ questions.

### v2

@@ 3.3 ADT-Auto Generation
Given the inefficiency of constructing dataset manually, we further develop an automatic generation framework for dataset to challenge LLMs’ tokenization. As discussed in Section 3.2, the challenges of tokenization in English are less severe than in Chinese. Consequently, we primarily concentrate on the automatic generation of Chinese data. Figure 3 illustrates the process of automatically constructing our dataset ADT-Auto.
@@ 3.3.1 Word-pair Matching
The automatic generation of dataset is also based on exported vocabularies. From the vocabularies, we first seek some qualified word pairs. Given two words Word 1 and Word 2, they are considered to be match when meeting the following criteria: The suffix (or whole) of Word 1 can be concatenated with the prefix (or whole) of Word 2, as a token existing in the vocabulary, denoted as Trap Word. Meanwhile, the concatenation of Word 1 and Word 2 should not be a token existing in the vocabulary.
@@
Above clarifications indicate that the concatenation of Word 1 and Word 2 (as ‘ABCD’ in Figure 3) corresponds to the challenging span in the construction process of ADT-Human. Our goal is to compose a sentence that not only convey the semantic essence of Word 1 and Word 2, but also induce LLM to isolate Trap Word when tokenization. Notably, we ignore the situation that the concatenation of Word 1 and Word 2 is just Trap Word, as it would not pose challenges to LLMs’ tokenization.
To augment the matching efficacy of Trap Word upon matching, we also consider the criterion that the remaining parts of Word 1 and Word 2 excluded by Trap Word should also exist in the vocabulary as a token. Furthermore, the first or last character of Trap Word cannot be a Chinese stop character.
@@ 3.3.2 Instance Generalization
With each Word 1, Word 2 and their corresponding Trap Word, we harness GPT-4 to generate an instance of ADT-Auto, which consists of a sentence and a question. Specifically, we prompt GPT-4 to generate a sentence that is used to challenge the tokenization of LLMs. To this end, the generated sentence is required to include the concatenation of Word 1 and Word 2 (inevitably including Trap Word). In addition, GPT-4 is also required to devise a question related to the generated sentence, which is used to evaluate LLMs’ tokenization performance through their answers. The instance must meet criteria as below: The sentence should convey the meanings of both Word 1 and Word 2, while the question’s answer should come from Word 1, Word 2, Trap Word or their combination. Thus, the influence of LLMs’ incorrect tokenization can be identified by checking answers to the question.
The prompt for GPT-4 to generate instances includes some demonstration examples in addition to task instruction, which is illustrated in Appendix D.
@@ 3.3.3 Filtering Based on Trap Word
The goal of our dataset is to expose the flaw of LLMs’ tokenization. Therefore, we will only retain the instances that can induce tokenization problem of LLMs, so we check each instance generated at the previous step. For these instances, the presence of Trap Word implies a challenging case that is likely to induce tokenization problems. Given an instance, we retain it if its corresponding Trap Word is both in the LLM’s tokenization list for the instance’s sentence (as List 1 in Figure 3) and in its tokenization list for the answer (as List 2 in Figure 3). Such filtering criterion indicates that the LLM commits tokenization errors on understanding the sentence and response for the instance.
@@ 3.3.4 Manual Filtering
To ensure the retained instances can induce tokenization problems and meanwhile have reasonable expressions, we further adopt manual assessment for instances. Specifically, we select the high-qualify instances with considering sentence expressions and LLM’s responses. Notably, we might still retain some instances to which the used LLM has correct response, since the other LLMs are still likely to commit inaccurate tokenization for these instances, resulting in unsatisfactory responses.
Due to space limitation, we take Qwen-7B as an example to illustrate the process of generating instances. There are 24,953 Chinese tokens in Qwen-7B’s vocabulary, and after word-pair matching, 1,764,692 word-pairs are obtained. From the matching word-pairs, 8,000 pairs are selected randomly and used for instance generation by GPT-4. Due to the inherent stochastic characteristic of LLMs on response generation, we conduct three iterations of filtering operations introduced in Subsection 3.3.3 to get more qualified instances, and thus retain 894 instances. Next, through the manual filtering introduced in Subsection 3.3.4, we retain 231 instances for ADT-Auto in the end.
@@ 4.1 Experiment Setup
Considering the open-source LLMs used in our experiments, we select the LLMs previously used in the construction of ADT-Human, including Chatglm3-6B, Baichuan2-13B-Chat, Yi-34B-Chat, Qwen-7B-Chat, and Qwen1.5-72B-Chat for Chinese data, as well as Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1 for English data. We test these LLMs using both locally deployed versions and API versions, with the exception of Chatglm3-6B, which does not have API version. For the closed-source LLMs, we test GPT-4o, GPT-4, GPT-3.5-Turbo, Qwen2.5-max, step-1-8k [^2] https://platform.stepfun.com/docs/llm/text, moonshot-v1-8k [^3] https://platform.moonshot.cn/docs, ERNIE-3.5-8K444 https://cloud.baidu.com/doc/WENXINWORKSHOP/s/jlil56u11 for Chinese data, and GPT-4o, GPT-4 and GPT-3.5-Turbo for English data. Additionally, we test the API version of Deepseek-R1, which has recently gained significant attention, for Chinese data.
@@ 4.2 ADT-Human’s Challenges to LLMs
Firstly, we investigate the challenges posed by the manually constructed dataset ADT-Human to LLMs. Specifically, we evaluate LLMs’ performance through counting the number of incorrect answers generated by LLMs for the questions in instances. Recalling the process of manually composing challenging spans introduced in Section 3.2, the span ‘BC’ in Table 2 is in fact the Trap Word mentioned in Section 3.3. Hence, for a given LLM, its answer including a Trap Word is identified as inaccurate response for the question undoubtedly. We identify the correctness of LLMs’ responses through human assessment.
The percentage of incorrect responses provided by the LLMs for Chinese data (corresponding to the four LLMs’ vocabularies) and English data (corresponding to the two LLMs’ vocabularies) are presented in Table 3 and Table 4, respectively. The results show that ADT-Human poses a significant challenge to both open-source and closed-source LLMs on their tokenization, resulting in very high rates of inaccurate responses. It is also worth mentioning that the recent state-of-the-art (SOTA) GPT-4o cannot outperform GPT-4 , implying that the advancements of these LLMs have not yet addressed this primary but challenging problem.
Model Source LLM of vocabulary Overall error rate Chatglm3 Baichuan2 Yi Qwen Open-source (Local) Chatglm3-6B 100.00 100.00 100.00 90.91 97.73 Baichuan2-13B-Chat 90.91 100.00 81.82 100.00 93.18 Yi-34B-Chat 72.73 63.64 100.00 100.00 84.09 Qwen-7B-Chat 100.00 72.73 90.91 100.00 90.91 Qwen1.5-72B-Chat 90.91 45.45 81.82 100.00 79.55 Open-source (API) Baichuan2-13B-Chat 100.00 100.00 90.91 100.00 97.73 Yi-34B-Chat 81.82 54.55 100.00 90.91 81.82 Qwen-7B-Chat 100.00 81.82 81.82 100.00 90.91 Qwen1.5-72B-Chat 90.91 54.55 81.82 100.00 81.82 DeepSeek-R1 36.36 27.27 45.45 54.55 40.91 Closed-source GPT-4o 72.73 27.27 54.55 90.91 61.36 GPT-4 81.82 45.45 45.45 27.28 50.00 GPT-3.5-Turbo 72.73 27.27 72.73 72.73 61.36 Qwen2.5-max 90.91 72.73 90.91 100.00 88.64 step-1-8k 63.64 18.18 72.73 63.64 54.55 moonshot-v1-8k 81.82 27.27 100.00 81.82 72.73 ERNIE-3.5-8K 72.73 54.55 54.55 72.73 63.64
@@
To further investigate the impact of tokenization on the performance of LLMs, a fine-grained analysis of the relationship between tokenization and response is conducted. In Appendix B, we quantitatively examine the relationship between the correctness of tokenization and the correctness of LLM’s response. Given the inaccessibility of tokenization lists for closed-source LLMs, we perform a statistical analysis on open-source LLMs. Since LLMs’ API versions do not directly provide tokenization lists, the tokenization results obtained from the corresponding locally deployed versions are used. Appendix B.1 and Appendix B.2 respectively illustrate the quantitative relationships between tokenization and response for the LLMs tested on Chinese and English data of ADT-Human.
Right Wrong
response response
Right tokenization 3 3
Wrong tokenization 1 37
Figure 4: Four relationships between tokenization and response, take Qwen-7B-Chat as an example.
This study further intuitively illustrates the relationships between tokenization and response in the form of pie charts in Appendix C. In Appendix C.1 and Appendix C.2, the proportions of four relationships between tokenization and response for each LLM on Chinese and English data of ADT-Human are displayed respectively. Figure 4 illustrates the quantitative relationships and pie chart using Qwen-7B-Chat as an example. For more details, please refer to Appendix B and Appendix C. We mainly focus on the proportion of TN (tokenization incorrect and response incorrect). As shown in pie charts, the proportion of TN cases is very high in ADT-Human, with an average of 80.91% for Chinese data and 79.78% for English data. This indicates that tokenization errors significantly affect the accuracy of LLM responses and also demonstrates that ADT-Human effectively challenges the performance of LLMs.
@@ 4.3 ADT-Auto’s Challenges to LLMs
Similar to the investigation of ADT-Human’s challenges to LLMs’ tokenization, we also evaluate the LLMs’ performance on ADT-Auto.
Model Fraction Error rate (%) Open-source (Local) Chatglm3-6B 156/231 67.53 Baichuan2-13B-Chat 147/231 63.64 Yi-34B-Chat 90/231 38.96 Qwen-7B-Chat 185/231 80.09 Qwen1.5-72B-Chat 93/231 40.26 Open-source (API) Baichuan2-13B-Chat 167/231 72.29 Yi-34B-Chat 80/231 34.63 Qwen-7B-Chat 160/231 69.26 Qwen1.5-72B-Chat 97/231 41.99 DeepSeek-R1 57/231 24.68 Closed-source GPT-4o 75/231 32.47 GPT-4 89/231 38.53 GPT-3.5-Turbo 99/231 42.86 Qwen2.5-max 75/231 32.47 step-1-8k 60/231 25.97 moonshot-v1-8k 60/231 25.97 ERNIE-3.5-8K 51/231 22.08
Table 5: Error rates of answers on ADT-Auto.
@@
    Compared with closed-source LLMs, open-source LLMs suffer from ADT-Auto’s challenges more apparently. It implies that these closed-source LLMs, as the profit-making flagships of their creator companies, naturally have stronger capabilities than open-source LLMs that are created for public usage.
@@
    Qwen1.5-72B-Chat has lower error rates than Qwen-7B-Chat, although they have the same vocabulary. We specially check some instances to which Qwen1.5-72B-Chat gives correct answers but Qwen-7B-Chat gives wrong answers, and find the two models have the same incorrect tokenization lists for these instances. It suggests that their different performance is not caused by tokenization. The results also imply that in the case of incorrect tokenization, the bigger LLMs are more robust and likely to generate correct responses than the smaller LLMs thanks to their stronger capabilities brought by the larger scale.
Similar to Section 4.2, Appendix B.3 presents the quantitative relationships between tokenization and response for each open-source LLM tested on ADT-Auto. The corresponding pie charts are shown in Appendix C.3. As indicated by the pie charts, the proportion of TN cases is also very high on ADT-Auto, with an average of 46.11%. This further demonstrates that tokenization errors significantly impact the accuracy of LLMs’ responses and highlights the effectiveness of ADT-Auto in challenging the performance of LLMs.
5 Conclusion
In this paper, we dedicate to deeply investigating the relationship between LLMs’ vulnerability on tokenization and their unsatisfactory responses for some tasks. To this end, we construct an adversarial dataset ADT (Adversarial Dataset for Tokenizer) consisting of a manually constructed subset ADT-Human and an automatically generated subset ADT-Auto, of which each instance includes a sentence and a corresponding question. Our experiments demonstrate that our dataset does challenge the studied open-source and closed-source LLMs’ token segmentation, resulting in their incorrect answers. We hope our work and dataset could shed light on the subsequent research on improving LLMs’ performance.
Limitations
The key contribution of this study lies in drawing attention to the impact of tokenization on LLM’s performance and providing two frameworks for data generation. As for how to propose improvement strategies based on this phenomenon, we are currently in the process of exploration. Additionally, this study focuses primarily on Chinese. Whether other languages are similarly affected by tokenization remains to be further investigated.
Ethics Statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations.
Human annotations are only used in methodological research at the beginning of the work, to assist in analyzing the feasibility of the proposed solution. Annotators consented to the use of data for research purposes. We ensure the privacy of all annotators is protected throughout the annotation process, and all of them are adequately paid according to local standards.
Risks.
Synthetic data generated by LLMs may involve potential ethical risks regarding fairness and bias Bommasani et al. (2021); Blodgett et al. (2020), which results in further consideration when they are employed in downstream tasks. We asked our members for proofreading to refine the offensive and harmful data generated by GPT-4. Despite these considerations, there may still be some unsatisfactory data that goes unnoticed in our final dataset.
@@ A.5 Llama-3
  1. 1.
    In the postal history course, today’s homework is to list|envelope|classes in history. What should I do?
  2. 2.
    As a Marathon lover, in the past he only ran|domestically, but now he also goes abroad. What behavior of him is discussed in the sentence?
  3. 3.
    As fans of Kenshi Yonezu, they sing|lemon in the auditorium. What song are they singing in the sentence?
  4. 4.
    I wanted to go to the beach, but|tons|of work piled up. How much work did I have?
  5. 5.
    We are measuring how fat|her|cat is to make sure her cat is healthy. What metric are we measuring?
  6. 6.
    This automotive shop mainly sells car|bonnets. What does the shop sell?
  7. 7.
    The researcher was disappointed to miss|ionization in the sample, which was crucial for the experiment’s success. What made the researcher disappointed?